# Márton Benedek 1122    Péter Biró 2211    Gergely Csáji 22 Matthew Johnson 33

### v2

@@ 4 Simulations
    to compare all these results with those for the other extreme case where $\ell=2$ [10].
@@ 4.1 The Set Up
To allow a fair comparison we use the same set up as in [10], and in addition, we also use the data from [6] that was used in [10]. That is, for our simulations, we take the same 100 compatibility graphs $G_{1},\ldots,G_{100}$, each with roughly 2000 vertices from [6, 10]. As real medical data is unavailable to us for privacy reasons, the data from [6] was obtained using the data generator from [34]. This data generator was used in many papers and is the most realistic synthetic data generator available; see also [24].
We use every $G_{i}$ ($i\in\{1,\ldots,100\}$) as the basis to perform simulations for $n$ countries, where we let $n=4,\ldots,10$. In order to do this, we first consider the case where all countries have equal size. That is, we partition $V(G_{i})$ into the same $n$ sets $V_{i,1},\ldots,V_{i,n}$ as in [10] which are of equal size ${2000}/{n}$ (subject to rounding), so $V_{i,p}$ is the set of patient-donor pairs of country $p$. For round 1, we construct a compatibility graph $G_{i}^{1}(n)$ as a subgraph of $G_{i}$ of size roughly $500$. We add the remaining patient-donor pairs of $G_{i}$ as vertices by a uniform distribution between the remaining rounds $2,\ldots,24$. So, starting with $G_{i}^{1}(n)$, we run an IKEP of 24 rounds in total. This gives us 24 compatibility graphs $G_{i}^{1}(n),\ldots,G_{i}^{24}(n)$. As in [10], any patient-donor pair whose patient is not helped within four rounds will automatically be deleted from the pool.
We also consider a situation where countries have different sizes. We note that there are numerous ways one could allow for differently sized countries. However, we follow the approach of [10], as this allows us to compare the results with those for $\ell=2$. That is, we consider the same 100 compatibility graphs $D_{1},\ldots,D_{100}$ as before; the only difference is that we now partition each $V(D_{i})$ into the same $n$ sets $V_{i,1},\ldots,V_{i,n}$ as in [10], such that
@@
Our code for obtaining the simulation instances can be found in a GitHub repository [44], along with the compatibility graphs data and the seeds for the randomization.
@@ 4.2 The Initial Allocations
We define each of these solution concepts below. Each of them prescribes exactly one allocation for the partitioned permutation games associated with the computability graphs. For computing these allocations we need the $v$-values of the partitioned permutation games.
We can compute a single value $v(S)$ in polynomial time by Theorem 1.1 through a transformation into a bipartite graph and solving a maximum weight perfect matching problem (see Section 1). For this we use the package LEMON version 1.3.1 of [25].
Let $(N,v)$ be a cooperative game. The Shapley value $\phi(N,v)$ [37] is defined by setting for $p\in N$,
@@
To define the next solution concept, we first introduce unnormalized Banzhaf value $\psi(N,v)$ [31] defined by setting for $p\in N$,
@@
We now define the nucleolus. The excess for an allocation $x$ of a game $(N,v)$ and a nonempty coalition $S\subsetneq N$ is defined as $e(S,x):=x(S)-v(S)$. Ordering the $22^{n}-$ 2 excesses in a non-decreasing sequence yields excess vector $e(x)\in\mathbb{R}^{2^{n}-2}$. The nucleolus of a game ($N,v$) is the unique allocation [36] that lexicographically maximizes $e(x)$ over the set of allocations $x$ with $x_{i}\geq v(\{i\})$ (assuming this set is nonempty, as is the case for partitioned permutation games). To compute the nucleolus, we use the Lexicographical Descent method of [11], which is the state-of-the-art method in nucleolus computation.
Next, we define the tau value, introduced by Tijs [41]. First, let $b_{p}=v(N)-v(N\setminus p)$ be the utopia payoff of $p\in N$, leading to a vector of utopia payoffs $b\in\mathbb{R}^{N}$. Now, for $p\in S\subseteq N$, we let
@@
The surplus of a game $(N,v)$ is defined as $\mbox{surp}=v(N)-\sum_{p\in N}v(\{p\})$. As long as $\sum_{p\in N}\alpha_{p}=1$, we can allocate $v(\{p\})+\alpha_{p}\cdot\mbox{surp}$ to each player $p\in N$. In this way we obtain the benefit value by setting for each $p\in N$,
$\alpha_{p}=\frac{v(N)-v(N\setminus\{p\})-v(\{p\})}{\sum_{p\in N}(v(N)-v(N\setminus\{p\})-v(\{p\}))},$
and we obtain the contribution value by setting for each $p\in N$,
$\alpha_{p}=\frac{v(N)-v(N\setminus\{p\})}{\sum_{p\in N}(v(N)-v(N\setminus\{p\}))}.$
@@
Both the benefit value and contribution value are easy to compute, as we only need to compute $12n-$ 1 $v$-values. We also note that that the benefit value and contribution value do not exist if the denominator is zero. However, this did not happen in our simulations.
@@
Finally, we consider a recent variant of the Banzhaf value, which was introduced in [10]. We first define for some round $h$ in an IKEP the credit-adjusted game $(N^{h},\overline{v}^{h})$, in which the credits are directly incorporated, that is, $\overline{v}^{h}(S)=v^{h}(S)+\sum_{p\in S}c^{h}_{p}$ for every $S\subseteq N$. As explained in [10], this alternative way of processing credits does not make any difference for any of the solution concepts that we consider except for one: the (normalized) Banzhaf value. We denote the adjusted Banzhaf value as the Banzhaf* value. In our simulations, we compute it by using brute force as we do for the Banzhaf value.
@@ 4.3 The Optimal Solutions
As optimal solutions we compute the following for each computability graph:
@@
In this way we can measure the effect of using weakly close optimal solutions over arbitrarily chosen ones as well as the effect of using strongly close optimal solutions over weakly close ones. We compute the weakly and strongly close optimal solutions by solving a sequence of ILPs, as described in Section 3. We use Gurobi Optimizer version 10.0.02(linux64) to create environments, build and solve ILPs (via the Gurobi C++ interface on a cloud system). As not every ILP might be solved within a reasonable time, we give each ILP a time limit of one hour. We recall that the values $v(S)$ of the partitioned permutation games used in the ILPs have already been computed for the initial allocations, and we simply reuse them.
@@ 4.4 Computational Environment and Scale
In our large-scale experimental study, we run our simulations both without and with (“$+c$”) the credit system, and as mentioned for the settings, where an arbitrary optimal solution (“arbitrary”), weakly optimal solution (“$d_{1}$”) or strongly close optimal solution (“lexmin”) is chosen. This leads to the following five scenarios for each of the six selected solution concepts, the Shapley value, Banzhaf value, nucleolus, tau value, benefit value and contribution value:
@@
as without credits the lhe Banzhaf* value coincides with the Banzhaf value.
Note that for the arbitrary scenario, the use of credits is irrelevant. Hence, in total, we run the same set of simulations for $5\times 6+2=32$ different combinations of scenarios and solution concepts. We consider both the situation where all countries have the same size and a situation where they have different sizes. Moreover, we have seven different country sizes $n$ and 100 compatibility graphs $G_{i}$. Hence, the scale of our experimental study is very large. Namely, the total number of 24-round simulation instances is
$32\times 2\times 7\times 100=44800$
All simulations were run on a dual socket server with AMD EPYC 7702 64-Core Processor with 2.00 GHz base speed and 256GB of RAM, where each simulation was given eight cores and 1G temporary disk space. As we will compare our results for $\ell=\infty$ with the known results for $\ell=2$ from [10], let us point out that the simulations in [10] were run on a desktop PC with AMD Ryzen 9 5950X 3.4 GHz CPU and 128 GB of RAM, running on Windows 10 OS and C++ implementation in Visual Studio (the code is available in GitHub repository [6]).
@@ 4.6 Simulation Results for ${\mathbf{\ell=\infty}}$ and Comparison with ${\mathbf{\ell=2}}$
In Figure 4 we display our main results for equal country sizes, that is, when all countries are of the same size (see also Section 4.1). In Figure 4, we compare different solution concepts under different scenarios for $\ell=\infty$. Figure 4 also shows the effects of weakly and strongly close solutions and the credit system. As solution concepts have different complexities, we believe such a comparison might be helpful for policy makers in choosing a specific solution concept and scenario. As expected, using an arbitrary maximum cycle packing in each round makes the kidney exchange scheme significantly more unbalanced, with average total relative deviations over 4% for all initial allocations $y$. The effect of both selecting a strongly close solution (to ensure being close to a target allocation) and using a credit function (for fairness, to keep deviations small) is significant.
@@
Figure 4: Average total relative deviations for each of the seven solution concepts under the five different scenarios for equal country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$. We recall that the Banzhaf value and Banzhaf* value coincide when credits are not incorporated, and this is also reflected in the two corresponding figures.
@@
The above observations are in line with the results under the setting where $\ell=2$ [10]. However, for $\ell=2$, the effect of using arbitrary optimal solutions is much worse, while deviations are smaller than for $\ell=\infty$ when weakly close or strongly close optimal solutions are chosen. From Figure 4 we see that the Banzhaf* value in the lexmin+c scenario provide the smallest deviations from the target allocations (as when $\ell=2$ [10]). However, all solution concepts are within 1.52% (for lexmin+c) and, as mentioned, which solution concept to select should be decided by the policy makers of the IKEP.
@@
Figure 5: Average total relative deviations for each of the seven solution concepts under the five different scenarios for varying country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$.
@@
Figure 5 shows the same kind of results as Figure 4 but now for varying country sizes, that is, when we have small, medium and large countries, divided exactly as in [10] for $\ell=2$ (see also Section 4.1). Note that subject to minor fluctuations we can draw the same conclusions from Figure 5 for varying country sizes as we did from Figure 4 for equal country sizes.
@@
Figures 4 and 5 highlight the comparison between different scenarios. For an easier comparison between the effects of choosing different solution concepts for prescribing the initial allocations, we grouped together all the lexmin+c plots in of Figures 4 and 5 in Figure 6 and all the d1+c plots from Figures 4 and 5 in Figure 7. Figures 6 and 7 both show an ordering from the Banzhaf* value, which has the best performance, to the contribution value which has the worst in almost all cases for the two most important scenarios lexmin+c and d1+c.
@@
We note that Figures 6 and 7 also show that (as expected) the effect of varying the country sizes is stronger if the number $n$ of countries is relatively small, especially when $n\in\{4,5,6\}$.
@@
Incomplete Instances. In the simulations for $\ell=2$ [10], ILPs were not used, and all simulation instances were solved to optimality. However, our new simulations for $\ell=\infty$ heavily relied on ILPs, and we recall that we imposed a time limit of one hour for our ILP solver (Gurobi) to solve an ILP. Given the nature of ILPs, it is not surprising that there were several ILPs that could not be solved within the $1$-hour time limit. We call the corresponding simulation instances incomplete. So, for such instances, there was one ILP in some round, which our ILP solver could not handle within the $1$-hour time limit. We call call other simulation instances complete.
All ILPs that were not finished within one hour were of the ($\text{ILP}_{d_{t}}$) type, almost always for $t>1$. In fact, for equal country sizes, only one single incomplete simulation instance occurred at d1 / d1+c (for nine countries, using the tau value). Tables 1 and 2 summarize the distribution of incomplete simulation instances aggregated over different choices of solution concepts or scenarios. We refer to Appendix 0.B for a complete breakdown of the averages in both these tables.
Given that the solutions found for incomplete simulation instances are all maximum cycle packings (which might not be weakly or strongly close), policy makers could therefore still decide to use them. For this reason, we decided to include the incomplete simulation instances in Figures 48. We believe this was justified after doing some additional research. That is, we also constructed the same figures as Figures 4 and 5 but without the incomplete simulation instances; see Appendix 0.B. It turned out that the largest percentage points difference in Figure 4 of the average total deviations is only 0.059% (with an average of 0.0024%). Hence, the quality of the “current-best” optimal solutions for the incomplete simulation instances are almost indistinguishably close to those for the complete simulation instances.
@@
Number of Kidney Transplants versus Cycle Length. When $\ell=\infty$ instead of $\ell=2$, we expect more kidney transplants as exchange cycles may now have any size. In Figures 8 and 9 we quantified this. From Figure 8 we see that when $\ell=\infty$ instead of $\ell=2$, it is possible to achieve 46% more kidney transplants. This is irrespective of the chosen scenario, as the total number of kidney transplants are nearly identical throughout the five scenarios (as for $\ell=2$). The $\ell=\infty$ setting is not realistic of course. Namely, Figure 9 show that long cycles, with even more than 400 vertices, may occur. Figure 9 also shows that these long cycles all happen in the first round. We note that before our experiments, the relation between increase in transplants versus increase in cycle length had not been researched by any simulations. Finding out about this relation through an extensive set of simulations was one of our main motivations.
Figure 8: Left: the total number of kidney transplants when the Banzhaf* value is chosen when the countries are of the same size. The dotted lines correspond to the case where $\ell=\infty$ and the solid lines are taken from [10] and correspond to the case where $\ell=2$. Right: the same figure but now for the case where the countries have different sizes. Note that in both figures some of the plots are overlapping.

### v3

@@ 4 Simulations
    to compare all these results with those for the other extreme case where $\ell=2$ [BBPY24].
@@ 4.1 The Set Up
To allow a fair comparison we use the same set up as in [BBPY24], and in addition, we also use the data from [Be21] that was used in [BBPY24]. That is, for our simulations, we take the same 100 compatibility graphs $G_{1},\ldots,G_{100}$, each with roughly [^2] The graphs $G_{i}$ might not have exactly 2000 vertices: depending on the specific number of countries that we consider, we may need to do some rounding in order for the countries to have the right sizes. 2000 vertices from [Be21, BBPY24]. As real medical data is unavailable to us for privacy reasons, the data from [Be21] was obtained using the data generator from [PT21]. This data generator was used in many papers and is the most realistic synthetic data generator available; see also [DGGKMPT22].
We use every $G_{i}$ ($i\in\{1,\ldots,100\}$) as the basis to perform simulations for $n$ countries, where we let $n=4,\ldots,10$. In order to do this, we first consider the case where all countries have equal size. That is, we partition the vertices $V_{i}$ of the graph $G_{i}$ into the same $n$ sets $V_{i,1},\ldots,V_{i,n}$ as in [BBPY24] which are of equal size 2000/n{2000}/{n} (subject to rounding), so $V_{i,p}$ is the set of patient-donor pairs of country $p$. For round 1, we construct a compatibility graph $G_{i}^{1}(n)$ as a subgraph of $G_{i}$ of size roughly $500$. We add the remaining patient-donor pairs of $G_{i}$ as vertices by a uniform distribution between the remaining rounds $2,\ldots,24$. So, starting with $G_{i}^{1}(n)$, we run an IKEP of 24 rounds in total. This gives us 24 compatibility graphs $G_{i}^{1}(n),\ldots,G_{i}^{24}(n)$. As in [BBPY24], any patient-donor pair whose patient is not helped within four rounds will automatically be deleted from the pool.
We also consider a situation where countries have different sizes. We note that there are numerous ways one could allow for differently sized countries. However, we follow the approach of [BBPY24], as this allows us to compare the results with those for $\ell=2$. That is, we consider the same 100 compatibility graphs $G_{1},\ldots,G_{100}$ as before; the only difference is that we now partition each $V_{i}$ into the same $n$ sets $V_{i,1},\ldots,V_{i,n}$ as in [BBPY24], such that
@@
Our code for obtaining the simulation instances can be found in a GitHub repository [Ye23], along with the compatibility graphs data and the seeds for the randomization.
@@ 4.2 The Initial Allocations
We define each of these solution concepts below. Each of them prescribes exactly one allocation for the partitioned permutation games associated with the compatibility graphs. For computing these allocations we need the $v$-values of the partitioned permutation games.
We can compute a single value $v(S)$ in polynomial time by Theorem 1.1 through a transformation into a bipartite graph and solving a maximum weight perfect matching problem (see Section 1). For this we use the package LEMON version 1.3.1 of [DJK11].
Let $(N,v)$ be a cooperative game. The Shapley value $\phi(N,v)$ [Sh53] is defined by setting for $p\in N$,
@@
To define the next solution concept, we first introduce unnormalized Banzhaf value $\psi(N,v)$ [Ba64] defined by setting for $p\in N$,
@@
We now define the nucleolus. The excess for an allocation $x$ of a game $(N,v)$ and a nonempty coalition $S\subsetneq N$ is defined as e(S,x):=x(S)v(S)e(S,x):=x(S)-v(S). Ordering the $2^{n}-2$ excesses in a non-decreasing sequence yields excess vector $e(x)\in\mathbb{R}^{2^{n}-2}$. The nucleolus $\eta$ of a game ($N,v$) is the unique allocation [Sc69] that lexicographically maximizes $e(x)$ over the set of allocations $x$ with $x_{i}\geq v(\{i\})$ (assuming this set is nonempty, as is the case for partitioned permutation games). To compute the nucleolus, we use the Lexicographical Descent method of [BFN21], which is the state-of-the-art method in nucleolus computation.
Next, we define the tau value, introduced by Tijs [Ti81]. First, let $b_{p}=v(N)-v(N\setminus p)$ be the utopia payoff of $p\in N$, leading to a vector of utopia payoffs $b\in\mathbb{R}^{N}$. Now, for $p\in S\subseteq N$, we let
@@
The surplus of a game $(N,v)$ is defined as $\mbox{surp}=v(N)-\sum_{p\in N}v(\{p\})$. As long as $\sum_{p\in N}\alpha_{p}=1$, we can allocate $v(\{p\})+\alpha_{p}\cdot\mbox{surp}$ to each player $p\in N$. In this way we obtain the benefit value [KNPV20] by setting for each $p\in N$,
$\alpha_{p}=\frac{v(N)-v(N\setminus\{p\})-v(\{p\})}{\sum_{{q}\in N}(v(N)-v(N\setminus\{{q}\})-v(\{{q}\}))},$
and we obtain the contribution value [BBKP22] by setting for each $p\in N$,
$\alpha_{p}=\frac{v(N)-v(N\setminus\{p\})}{\sum_{{q}\in N}(v(N)-v(N\setminus\{{q}\}))}.$
@@
Both the benefit value and contribution value are easy to compute, as we only need to compute 2n12n-1 $v$-values. We also note that that the benefit value and contribution value do not exist if the denominator is zero. However, this did not happen in our simulations.
@@
Finally, we consider a recent variant of the Banzhaf value, which was introduced in [BBPY24]. We first define for some round $h$ in an IKEP the credit-adjusted game $(N^{h},\overline{v}^{h})$, in which the credits are directly incorporated, that is, $\overline{v}^{h}(S)=v^{h}(S)+\sum_{p\in S}c^{h}_{p}$ for every $S\subseteq N$. As explained in [BBPY24], this alternative way of processing credits does not make any difference for any of the solution concepts that we consider except for one: the (normalized) Banzhaf value. We denote the adjusted Banzhaf value as the Banzhaf* value. In our simulations, we compute it by using brute force as we do for the Banzhaf value.
The solution concepts defined above have several fairness properties. For example, they may prescribe an allocation $x_{i}$ for a game $(N,v)$ that is individual rational, that is, $x_{p}\geq v(\{p\})$ for every $p\in N$. This property holds for the nucleolus, benefit value and contribution value by definition. For the tau value, which is only defined on quasibalanced games, it follows from the observation that for every $p\in N$, $\tau_{p}\geq a_{p}\geq v(\{p\})$. The Shapley value and unnormalized Banzhaf value are not individual rational in general. However, they are readily seen to be individual rational for superadditive games, i.e., games $(N,v)$ for which $v(S\cup T)\geq v(S)+v(T)$ for all pairs $S,T\subset N$ with $S\cap T=\emptyset$, which include the partitioned permutation games. In contrast, the normalized Banzhaf value is not individual rational for partitioned permutation games, as the following example shows.
Example 3. Let $G=(\{a,b,c\},\{(a,b),(b,a),(b,c),(c,b)\}$ and $V_{1}=\{a\}$, $V_{2}=\{b\}$ and $V_{3}=\{c\}$. Now, $\eta=(0,2,0)$ is the only core allocation, and we also note that $\phi_{1},\psi_{1},\overline{\psi}_{1}>0$. If we extend $G$ by adding $V_{4}=\{d,e\}$ with arcs $(d,e)$, $(e,d)$, then $\psi_{4}=v(\{4\})=2>\overline{\psi}_{4}$. ∎
As discussed in Section 1, the nucleolus $\eta$ is not only individual rational but even belongs to the core (so $\eta(S)\geq v(S)$ for every $S\subseteq N$). For a more detailed description on these solution concepts and their fairness properties, we refer to [Peters2008, Young1985]. Here, we only highlight one other fairness property to illustrate that policy makers must make a choice on what fairness properties they find most important (in addition to any computational considerations). A solution concept for a game $(N,v)$ is coalitional monotone if for every $i\in N$ the following holds: if the value of each of the coalitions player $i$ is a member of weakly increases, while the values of all the other coalitions remain the same, then the payoff of player $i$ must also weakly increase. The Shapley value and the Banzhaf value have this fairness property, whilst the nucleolus does not [Young1985]. However, whereas the nucleolus always belongs to the core of the game, Example 3 shows that the Shapley and Banzhaf values of permutation games may not be core allocations. In fact, Young [Young1985] even proved that no coalition monotone solution concept can guarantee to prescribe a core allocation.
Fairness properties, such as individual rationality, core membership and coalition monotonicity, are highly relevant for decision and policy making in international kidney exchange as well. As we have seen, such properties may be in conflict with each other. Hence, IKEPs should decide which fairness properties they prefer when selecting a certain solution concept. We also note that (fractional) target allocations are typically not possible to achieve exactly. This is, for instance, illustrated by Example 3, where for the original compatibility graph $G$, the nucleolus is $(0,2,0)$ while the only two optimal solutions are $(1,1,0)$ and $(0,1,1)$. However, the credit system offers a way to compensate for any deviations from target allocations in the long term.
@@ 4.3 The Optimal Solutions
As optimal solutions we compute the following for each compatibility graph:
@@
In this way we can measure the effect of using weakly close optimal solutions over arbitrarily chosen ones as well as the effect of using strongly close optimal solutions over weakly close ones. We compute the weakly and strongly close optimal solutions by solving a sequence of ILPs, as described in Section 3. We use Gurobi Optimizer version 10.0.02(linux64) to create environments, build and solve ILPs (via the Gurobi C++ interface on a cloud system). As not every ILP might be solved within a reasonable time, we give each ILP a time limit of one hour. As a minor comment, we recall that the values $v(N)$ of the partitioned permutation games used in the ILPs have already been computed for the initial allocations, and we simply reuse these to save some computation time.
@@ 4.4 Computational Environment and Scale
In our large-scale experimental study, we run our simulations both without and with (“+c+c”) the credit system, and as mentioned for the settings, where an arbitrary optimal solution (“arbitrary”), weakly optimal solution ($d_{1}$”) or strongly close optimal solution (“lexmin”) is chosen. This leads to the following five scenarios for each of the six selected solution concepts, the Shapley value, Banzhaf value, nucleolus, tau value, benefit value and contribution value:
@@
as without credits the Banzhaf* value coincides with the Banzhaf value.
Note that for the arbitrary scenario, the use of credits is irrelevant. Hence, in total, we run the same set of simulations for $5\times 6+2=32$ different combinations of scenarios and solution concepts. We consider both the situation where all countries have the same size and a situation where they have different sizes. Moreover, we vary the number of countries $n$ from $4$ to $10$ (so seven different numbers) and have 100 compatibility graphs $G_{i}$. Hence, the scale of our experimental study is very large. Namely, the total number of 24-round simulation instances is
$32\times 2\times 7\times 100=44800.$
All simulations were run on a dual socket server with AMD EPYC 7702 64-Core Processor with 2.00 GHz base speed and 256GB of RAM, where each simulation was given eight cores and 1G temporary disk space. As we will compare our results for $\ell=\infty$ with the known results for $\ell=2$ from [BBPY24], let us point out that the simulations in [BBPY24] were run on a desktop PC with AMD Ryzen 9 5950X 3.4 GHz CPU and 128 GB of RAM, running on Windows 10 OS and C++ implementation in Visual Studio (the code is available in GitHub repository [Be21]).
@@ 4.6 Simulation Results for $\mathbf{\ell=\infty}$} and Comparison with $\mathbf{\ell=2}$}
In this section, we present our simulation results for $\ell=\infty$ and compare these results with the simulation results for $\ell=2$ [BBPY24]. We start with Figure 5. In this figure we display our main results for equal country sizes, that is, when all countries are of the same size (see also Section 4.1), and we compare the seven different solution concepts under the five different scenarios for $\ell=\infty$ with each other. Figure 5 also shows the effects of weakly and strongly close solutions and the credit system. As solution concepts have different complexities, we believe such a comparison might be helpful for policy makers in choosing a specific solution concept and scenario. Figure 5 shows that using an arbitrary maximum cycle packing in each round indeed makes the kidney exchange scheme significantly more unbalanced, with average total relative deviations over 4% for all initial allocations $y$. Figure 5 also shows that the effect of both selecting a strongly close solution (to ensure being close to a target allocation) and using a credit function (for fairness, to keep deviations small) is significant. Finally, we observe that for all solution concepts and all scenarios, the deviations are slowly increasing when the number of countries is increasing. However, the rate of increase is slower for strongly close optimal solutions than for weakly close optimal solutions, while the “starting” point of the line is lower when credits are employed.
@@
Figure 5: Average total relative deviations for each of the seven solution concepts under the five different scenarios for equal country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [BBPY24] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$. We recall that the Banzhaf value and Banzhaf* value coincide when credits are not incorporated, and this is also reflected in the two corresponding figures.
@@
The above conclusions are in line with the results under the setting where $\ell=2$ [BBPY24]. However, for $\ell=2$, the effect of using arbitrary optimal solutions is much worse, while deviations are smaller than for $\ell=\infty$ when weakly close or strongly close optimal solutions are chosen. In Figure 5 we illustrated this for the Banzhaf* value. From Figure 5 we see that the Banzhaf* value in the lexmin+c scenario provides the smallest deviations from the target allocations (as when $\ell=2$ [BBPY24]). However, all solution concepts are within 1.52% (for lexmin+c) and, as mentioned, which solution concept to select should be decided by the policy makers of the IKEP.
@@
Figure 6: Average total relative deviations for each of the seven solution concepts under the five different scenarios for varying country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [BBPY24] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$.
@@
Figure 6 shows the same kind of results as Figure 5 but now for varying country sizes, that is, when we have small, medium and large countries, divided exactly as in [BBPY24] for $\ell=2$ (see also Section 4.1). Note that subject to minor fluctuations we can draw the same conclusions from Figure 6 for varying country sizes as we did from Figure 5 for equal country sizes.
@@
Figures 5 and 6 highlight the comparison between different scenarios. For an easier comparison between the effects of choosing different solution concepts for prescribing the initial allocations, we grouped together all the lexmin+c plots in of Figures 5 and 6 in Figure 7 and all the d1+c plots from Figures 5 and 6 in Figure 8. Figures 7 and 8 both show an ordering from the Banzhaf* value, which has the best performance, to the contribution value which has the worst in almost all cases for the two most important scenarios lexmin+c and d1+c.
@@
We note that Figures 7 and 8 also show that (as expected) the effect of varying the country sizes is stronger if the number $n$ of countries is relatively small, especially when $n\in\{4,5,6\}$.
@@
Incomplete Instances. In the simulations for $\ell=2$ [BBPY24], ILPs were not used, and all simulation instances were solved to optimality. However, our new simulations for $\ell=\infty$ heavily relied on ILPs, and we recall that we imposed a time limit of one hour for our ILP solver (Gurobi) to solve an ILP. Given the nature of ILPs, it is not surprising that there were several ILPs that could not be solved within the $1$-hour time limit. We call the corresponding simulation instances incomplete. So, for such instances, there was one ILP in some round, which our ILP solver could not handle within the $1$-hour time limit. We call call other simulation instances complete.
All ILPs that were not finished within one hour were of the ($\text{ILP}_{d_{t}}$) type, almost always for $t>1$. In fact, for equal country sizes, only one single incomplete simulation instance occurred at d1 / d1+c (for nine countries, using the tau value). Tables 2 and 3 summarize the distribution of incomplete simulation instances aggregated over different choices of solution concepts or scenarios. See Appendix 0.B for a complete breakdown of the averages in both these tables. Here, we only observe from Tables 2 and 3 that the average number of incomplete instances is relatively low, both for equal country sizes (at most 2% in total) and varying country sizes (at most 0.75% in total).
Given that the average number of incomplete instances is relatively low, and moreover that the solutions found for incomplete simulation instances are all maximum cycle packings (which might not be weakly or strongly close), policy makers could therefore still decide to use them. For this reason, we decided to include the incomplete simulation instances in Figures 58. We believe this was justified after doing some additional research. That is, we also constructed the same figures as Figures 5 and 6 but without the incomplete simulation instances; see Appendix 0.B. It turned out that the largest percentage points difference in Figure 5 of the average total deviations is only 0.059% (with an average of 0.0024%). Hence, the quality of the “current-best” optimal solutions for the incomplete simulation instances are almost indistinguishably close to those for the complete simulation instances.
@@
Number of Kidney Transplants and Cycle Length. When $\ell=\infty$ instead of $\ell=2$, we expect more kidney transplants, because exchange cycles may now have any size. In Tables 4-5 and Figure 9 we quantified this.
First, from Table 4 we see that for $n=10$ and equal country sizes it is possible to achieve 46% more kidney transplants when $\ell=\infty$ instead of $\ell=2$, and that this is irrespective of the chosen scenario or chosen solution concept, as the total number of kidney transplants are nearly identical, both for $\ell=2$ and $\ell=\infty$. Table 5 shows the results for varying country sizes, for which we can draw the same conclusions. We refer to Tables 19 and 20 in Appendix 0.D for the corresponding results for $n=4,\ldots,9$, which show the same behaviour.
Second, Figure 9 shows that setting $\ell=\infty$ may indeed lead to long cycles, with even more than 400 vertices. Figure 9 shows in particular that these long cycles all occur in the first round. The reason is that we start the first round with roughly 500 vertices and introduce the remaining 1500 patient-donor pairs in the later rounds by a uniform distribution. Hence, the compatibility graphs in later rounds are all smaller.
Solution concepts/scenarios arbitrary d1 d1+c lemxin lexmin+c
$\ell=\infty$ benefit value 1781.58 1781.65 1781.46 1782.58 1780.31
contribution value 1781.58 1781.52 1780.50 1782.52 1780.05
Nucleolus 1781.58 1781.80 1780.45 1781.98 1780.12
Shapley value 1781.58 1781.54 1780.92 1782.02 1780.57
Banzhaf value 1781.58 1781.83 1780.38 1782.46 1780.66
Banzhaf* value 1781.58 1781.83 1781.08 1782.46 1781.24
Tau value 1781.58 1782.22 1780.90 1782.12 1780.45
$\ell=2$ benefit value 1221.38 1222.4 1223.44 1222.54 1223.6
contribution value 1221.38 1222.58 1222.56 1223.24 1222.66
Nucleolus 1221.38 1222.28 1221.18 1221.84 1221.66
Shapley value 1221.38 1221.6 1221.2 1223.04 1222.32
Banzhaf value 1221.38 1222.62 1221.26 1220.9 1221.98
Banzhaf* value 1221.38 1222.62 1220.86 1220.9 1221.82
Tau value 1221.38 1223.96 1223.24 1224.26 1222.72
Table 4: Average number of kidney transplants for equal country sizes for $n=10$ when $\ell=2$ [BBPY24] and $\ell=\infty$.
Solution concepts/scenarios arbitrary d1 d1+c lemxin lexmin+c
$\ell=\infty$ benefit value 1763.57 1763.39 1763.37 1763.73 1762.45
contribution value 1763.57 1763.45 1762.90 1763.80 1761.84
Nucleolus 1763.66 1764.02 1762.79 1763.74 1762.19
Shapley value 1764.04 1763.09 1762.99 1763.71 1762.60
Banzhaf value 1763.24 1763.34 1762.85 1763.47 1762.85
Banzhaf* value 1763.24 1763.34 1763.09 1763.47 1762.97
Tau value 1763.27 1763.54 1762.99 1763.62 1761.84
$\ell=2$ benefit value 1203.8 1206.14 1205.92 1205.06 1205.88
contribution value 1204.06 1205.7 1205.46 1205.36 1206.16
Nucleolus 1204.06 1205.34 1205.38 1205.52 1207
Shapley value 1204.06 1205.7 1206.64 1206.34 1205.48
Banzhaf value 1204.06 1206.14 1205.56 1205.88 1206.34
Banzhaf* value 1204.06 1206.14 1206.34 1205.88 1205.92
Tau value 1204.06 1206.86 1205.92 1207.32 1206.3
Table 5: Average number of kidney transplants for varying country sizes for $n=10$ when $\ell=2$ [BBPY24] and $\ell=\infty$.